# Infinite graph product of groups I: Geometry of the extension graph

### v2

@@ Abstract
We introduce the extension graph of graph product of groups and study its geometry. This enables us to study properties of graph product by exploiting large scale geometry of its defining graph. In particular, we show that the extension graph is isomorphic to the crossing graph of a canonical quasi-median graph and exhibits the same phenomenon about asymptotic dimension as quasi-trees of metric spaces studied by Bestvina-Bromberg-Fujiwara. As an application of the extension graph, we prove relative hyperbolicity of graph-wreath product. This provides a new construction of relatively hyperbolic groups.
@@ 1 Introduction
However, most results so far concern only the case where the defining graph $\Gamma$ is finite, or the results just hold for any defining graph. To the best of my knowledge, there is no systematic study about how the geometry of a defining graph affects properties of graph product of groups and the purpose of this paper is to investigate this direction of research. In order to capture the geometry of a defining graph, it is not effective to consider graph product as amalgamated free product or to consider Cayley graphs, which are standard methods so far to study graph product. Instead, we achieve it by constructing a graph $\Gamma^{e}$ that remembers the geometry of the defining graph $\Gamma$ and also admits an action by the graph product. This graph $\Gamma^{e}$, which we call the extension graph (see Definition 3.1), generalizes the extension graph constructed for right angled Artin groups (RAAGs) by Kim and Koberda (see [KK13] and [KK14]).
@@
However, there is a crucial difference in the general case of graph product. That is, the extension graph $\Gamma^{e}$ in general case cannot be obtained by repeatedly taking the double of a graph along a star unlike the case of RAAGs. More precisely, [KK13, Lemma 22], which is a key lemma in [KK13], is no longer true in general (see Lemma 3.21). Therefore, we have to develop a new way to study the extension graph of general graph product, which we do in Section 3. For a technical reason, we have to assume that the girth of a defining graph is more than 20 in many of our results. Although this condition on the girth may well not be optimal, it is often essential in our applications. Also, we can increase the girth of a graph easily by taking its barycentric subdivision.
@@
Moreover, it turns out that our extension graph $\Gamma^{e}$ and its coned-off graph $\widehat{\Gamma}^{e}$ (see Definition 3.17) are isomorphic to the crossing graph and the contact graph respectively of a canonical quasi-median graph associated to graph product (see Section 3.2). This generalizes observation by Genevois in the case of Kim-Koberda’s extension graph of RAAGs (see [Gen17, Section 8.5]). Therefore, this paper provides a new perspective to study the crossing graph and the contact graph.
@@
One feature of the Kim-Koberda’s extension graph of right angled Artin groups was that it is quasi-isometric to a tree when the defining graph is connected (see [KK13, Lemma 26 (7)]). This is no longer true in the general case of graph product, because the defining graph $\Gamma$, which we assume is connected throughout this paper, isometrically embeds into the extension graph $\Gamma^{e}$ (see Corollary 3.30). However, this fact for RAAGs means that when the asymptotic dimension of a defining graph is 0, the asymptotic dimension of the extension graph is at most 1. We generalize this as follows.
@@
In this paper, we present one application of Theorem 1.2 to graph-wreath product. In the forthcoming paper, we present more applications to analytic properties of graph product. Given a group $G$ acting on a simplicial graph $\Gamma$ and another group $H$, we can assign vertex groups $\mathcal{G}=\{G_{v}\}_{v\in V(\Gamma)}$ by setting $G_{v}=H$ for every $v\in V(\Gamma)$, and the group $G$ acts on the graph product $\Gamma\mathcal{G}$ as group automorphisms by permuting vertex groups according to the action $G\curvearrowright\Gamma$. This action $G\to\mathrm{Aut}(\Gamma\mathcal{G})$ defines the semi-direct product $\Gamma\mathcal{G}\rtimes G$, which was called graph-wreath product in [KM16]. Interestingly, we can consider this construction as interpolation of wreath product and free product. Indeed, given groups $G$ and $H$, the group $G$ acts on the set $G$ by left multiplication. If $\Gamma$ is the complete graph with $V(\Gamma)=G$, then we have $\Gamma\mathcal{G}\rtimes G=H\wr G$. If $\Gamma$ is the graph with $V(\Gamma)=G$ having no edge, then we have $\Gamma\mathcal{G}\rtimes G=H\ast G\,(=\ast_{g\in G}gHg^{-1}\rtimes G)$.
In [KM16], Kropholler and Martino studied homotopical finiteness conditions, i.e. type $\mathrm{F}_{n}$, of graph-wreath product. In Theorem 1.3 and Corollary 1.4 below, we prove geometric properties of graph-wreath product, which provide a new construction of relatively hyperbolic groups and hyperbolic groups.
@@
    For any $v\in V(G)$, $\mathrm{Stab}_{G}(v)$ is finitely generated.
@@
It seems interesting to investigate properties of hyperbolic groups constructed in Corollary 1.4. For example, studying residual finiteness of these hyperbolic groups might shed light on the long-standing open problem asking whether every hyperbolic group is residually finite, because non residually finite groups can be easily constructed by wreath product. Indeed, wreath product of two groups $H\wr G$ is residually finite if and only if $H$ and $G$ are residually finite, and either $H$ is abelian or $G$ is finite (see [Gru57, Theorem 3.2]).
This paper is organized as follows. In Section 2, we explain preliminary definitions and known results that are necessary in this paper. In Section 3, we define the extension graph of graph product and study its geometry. In Section 4, we study asymptotic dimension of the extension graph and prove Theorem 1.1 by introducing an auxiliary graph that we call the coned-off extension graph. In Section 5, we discuss how properties of a defining graph carry over to the extension graph and prove Theorem 1.2 by using results in Section 3. In Section 6, we discusses the application to graph-wreath product and prove Theorem 1.3 and Corollary 1.4.
Acknowledgment. I would like to thank Anthony Genevois and Denis Osin for helpful discussions. I would also like to thank Peter Kropholler for helpful comments on an earlier draft.
@@ 2 Preliminaries
$\mathrm{Lk}_{X}(x)$ =$=\{y\in V(X)\mid(x,y)\in E(X)\},$
$\mathrm{St}_{X}(x)$ =$=\{x\}\cup\mathrm{Lk}_{X}(x),$
$\mathrm{Elk}_{X}(x)$ =$=\{(x,y)\in E(X)\mid y\in\mathrm{Lk}_{X}(x)\}.$
@@
Definition 2.6.
Let $(X,d_{X})$ be a metric space. For a subset $A\subset X$, the diameter of $A$ $\mathrm{diam}_{X}(A)\in[0,\infty]$ is defined by $\mathrm{diam}_{X}(A)=\sup_{x,y\in A}d_{X}(x,y)$. For $A\subset X$ and $r\in\mathbb{R}_{>0}$, we define $\mathcal{N}_{X}(A,r)\subset X$ by $\mathcal{N}_{X}(A,r)=\{y\in X\mid\exists\,x\in A,d_{X}(x,y)\leq r\}$. When $A$ is a singleton i.e. $A=\{x\}$ with $x\in X$, we denote $\mathcal{N}_{X}(\{x\},r)$ by $\mathcal{N}(x,r)$ for brevity, that is, $\mathcal{N}(x,r)=\{y\in X\mid d_{X}(x,y)\leq r\}$. For two subsets $A,B\subset X$, we define $d_{X}(A,B)\in\mathbb{R}_{\geq 0}$ by $d_{X}(A,B)=\inf_{x\in A,y\in B}d_{X}(x,y)$.
Definition 2.7.
Let $G$ be a group. For $g,h\in G$, we define $[g,h]\in G$, by $[g,h]=ghg^{-1}h^{-1}$. For subsets $A,B\subset G$, we define $AB,[A,B]\subset G$ by $AB=\{gh\in G\mid g\in A,h\in B\}$ and $[A,B]=\{[g,h]\in G\mid g\in A,h\in B\}$. For a subset $A\subset G$, we denote by $\langle A\rangle$ the subgroup of $G$ generated by $A$ and also by $\langle\!\langle A\rangle\!\rangle$ the normal subgroup of $G$ generated by $A$, that is, $\langle\!\langle A\rangle\!\rangle=\langle\bigcup_{g\in G}gAg^{-1}\rangle$.
Definition 2.8.
@@ 2.1 Graph products of groups
Let $\Gamma$ be a simplicial graph and $\mathcal{G}=\{G_{v}\}_{v\in V(\Gamma)}$ be a collection of groups. For any $g\in\Gamma\mathcal{G}$ with $g\neq 1$, $g=g_{1}\cdots g_{n}$ is a normal form of $g$ if and only if for any pair $(i,j)$ with $1\leq i<j\leq n$ satisfying $v_{i}=v_{j}$, there exists $k$ with $i<k<j$ such that $(v_{k},v_{i})\notin E(\Gamma)$. Also, if $g=g_{1}\cdots g_{n}$ and $g=h_{1}\cdots h_{m}$ are normal forms of $g$, then $n=m$ and we obtain one from the other by finite steps of syllable shuffling.
@@ 2.2 The crossing graph and the contact graph of a quasi-median graph
Let $X$ be a quasi-median graph. A hyperplane is an equivalence class of edges of $X$, where two edges $e$ and $f$ of $X$ are defined to be equivalent if there exists a sequence of edges $e=e_{0},\cdots,e_{n}=f$ of $X$ such that, for every $11\leq i\leq n-$ 1, $e_{i}$ and $e_{i+1}$ are either two sides of a triangle or opposite sides of a square. We denote by $\mathcal{H}(X)$ the set of all hyperplanes of $X$. For a hyperplane $J$ of $X$, we denote by $N^{\prime}(J)$ the set of the endpoints of all the edges in $J$. A carrier of $J\in\mathcal{H}(X)$, denoted by $N(J)$, is the induced subgraph on $N^{\prime}(J)$. The crossing graph of $X$, denoted by $\Delta X$, is a graph with $V(\Delta X)=\mathcal{H}(X)$, where two distinct vertices $J_{1},J_{2}\in\mathcal{H}(X)$ are adjacent if there exists a square in $X$ that contains both an edge in $J_{1}$ and an edge in $J_{2}$. The contact graph of $X$, denoted by $\mathcal{C}X$, is a graph with $V(\mathcal{C}X)=\mathcal{H}(X)$, where two distinct vertices $J_{1},J_{2}\in\mathcal{H}(X)$ are adjacent if $N^{\prime}(J_{1})\cap N^{\prime}(J_{2})\neq\emptyset$.
@@ 3.1 Definition and basic properties of the extension graph of graph product
$V(\Gamma^{e})$ =$=\{gG_{v}g^{-1}\in 2^{\Gamma\mathcal{G}}\mid v\in V(\Gamma),g\in\Gamma\mathcal{G}\},$
$E(\Gamma^{e})$ =$=\{(gG_{v}g^{-1},hG_{w}h^{-1})\in V(\Gamma^{e})^{2}\mid\text{$gG_{v}g^{-1}\neq hG_{w}h^{-1}$ and $[gG_{v}g^{-1},hG_{w}h^{-1}]=\{1\}$}\}.$
@@
We can assume $h=1$ without loss of generality, because we can consider $(h^{-1}g.G_{v},G_{w})\in E(\Gamma^{e})$ in general case. Since $G_{v}$ and $G_{w}$ are non-trivial and we have $(g.G_{v},G_{w})\in E(\Gamma)$, there exist $a\in G_{v}\setminus\{1\}$ and $b\in G_{w}\setminus\{1\}$ such that $[gag^{-1},b]=1$. By $(gag^{-1})b(gag^{-1})^{-1}=b$ and Corollary 3.5, we have $\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)$. This and Lemma 3.4 imply $v\in\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)$. Suppose $v=w$ for contradiction, then we have $\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)=\mathrm{St}_{\Gamma}(v)$. This implies $\mathrm{supp}(g)\subset\mathrm{St}_{\Gamma}(v)$ by Lemma 3.3 and Lemma 3.4. Hence, $g.G_{v}=G_{v}=G_{w}$. This contradicts $g.G_{v}\neq G_{w}$ by $(g.G_{v},G_{w})\in E(\Gamma^{e})$. Hence, $v\neq w$. By this and $v\in\mathrm{St}_{\Gamma}(w)$, we have $(v,w)\in E(\Gamma)$. By Lemma 3.4, there exist $h_{1},h_{2},h_{3}\in\Gamma\mathcal{G}$ such that $g=h_{1}h_{2}h_{3}$, $\mathrm{supp}(h_{1})\cup\{v\}=\mathrm{supp}(gag^{-1})$, $\mathrm{supp}(h_{2})\subset\mathrm{Lk}_{\Gamma}(v)$, and $\mathrm{supp}(h_{3})\subset\{v\}$. By $\mathrm{supp}(h_{2}h_{3})\subset\mathrm{supp}(h_{2})\cup\mathrm{supp}(h_{3})\subset\mathrm{St}_{\Gamma}(v)$, we have $g.G_{v}=h_{1}.G_{v}$. On the other hand, by $\mathrm{supp}(h_{1})\subset\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)$, we have $G_{w}=h_{1}.G_{w}$. ∎
@@
[…]
@@
    For any $a,b\in V(\Gamma^{e})$ with $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\neq\{1\}$, we have $d_{\Gamma^{e}}(a,b)\leq 4$.
@@
    If in addition $\mathrm{girth}(\Gamma)>5$ and the collection $\mathcal{G}=\{G_{v}\}_{v\in V(\Gamma)}$ is non-trivial finite groups, then for any $v,w\in V(\Gamma)$ with $v\neq w$ and any $g\in\Gamma\mathcal{G}$, the group $\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}$ is finite.

### v3

@@ Abstract
We introduce the extension graph of graph product of groups and study its geometry. This enables us to study properties of graph product by exploiting large scale geometry of its defining graph. In particular, we show that the extension graph is isomorphic to the crossing graph of a canonical quasi-median graph and exhibits the same phenomenon about asymptotic dimension as quasi-trees of metric spaces studied by Bestvina-Bromberg-Fujiwara. As an application of the extension graph, we prove relative hyperbolicity of graph wreath product. This provides a new construction of relatively hyperbolic groups.
Contents
@@ 1 Introduction
MSC 2020 Primary: 20F65. Secondary: 20F67, 51F30.Key words and phrases: graph product, asymptotic dimension, relatively hyperbolic groups.
@@
However, most results so far concern only the case where the defining graph $\Gamma$ is finite, or the results just hold for any defining graph. To the best of my knowledge, there is no systematic study about how the geometry of a defining graph affects properties of graph product of groups and the purpose of this paper is to investigate this direction of research. In order to capture the geometry of a defining graph, it is not effective to consider graph product as amalgamated free product or to consider Cayley graphs, which are standard methods so far to study graph product. Instead, we achieve it by studying a graph $\Gamma^{e}$ that remembers the geometry of the defining graph $\Gamma$ and also admits an action by the graph product. This graph $\Gamma^{e}$, which we call the extension graph (see Definition 3.1), was introduced by various people (see [Gen17, CRKdlNG24, EH24]) and generalizes the extension graph constructed for right angled Artin groups (RAAGs) by Kim and Koberda (see [KK13] and [KK14]).
@@
However, there is a crucial difference in the general case of graph product from that of RAAGs. That is, the extension graph $\Gamma^{e}$ in general case cannot be obtained by repeatedly taking the double of a graph along a star unlike the case of RAAGs. More precisely, [KK13, Lemma 22], which is a key lemma in [KK13], is no longer true in general (see Lemma 5.11). Therefore, we develop a new way to study the extension graph of general graph product, which we do in Section 3. For a technical reason, we have to assume that the girth of a defining graph is more than 20 in many of our results. Although this condition on the girth may well not be optimal, it is often essential in our applications. Also, we can increase the girth of a graph easily by taking its barycentric subdivision and assigning vertex groups to the new vertices. For various groups with features of non-positive curvature, we can also make a Cayley graph have arbitrarily large girth by choosing a nice finite generating set (see [Akh05, Yam11, Nak14, Akh25]).
@@
Moreover, it turns out that our extension graph $\Gamma^{e}$ and its coned-off graph $\widehat{\Gamma}^{e}$ (see Definition 3.17) are isomorphic to the crossing graph and the contact graph respectively of a canonical quasi-median graph associated to graph product (see Section 3.2). This generalizes the observation of Kim and Koberda in the case of the extension graph of RAAGs (see [KK14, Section 7], [Gen17, Section 8.5]). Therefore, this paper provides a new perspective to study the crossing graph and the contact graph by using the geometry of a defining graph $\Gamma$. (Another way to study these graphs is to use the geometry of a quasi-median Cayley graph of graph product groups (see [Gen17, Val21]).)
@@
One feature of the Kim-Koberda’s extension graph of right angled Artin groups was that it is quasi-isometric to a tree when the defining graph is connected (see [KK13, Lemma 26 (7)]). This is no longer true for general graph product once we allow infinite defining graphs. It is because the defining graph $\Gamma$, which we assume is connected throughout this paper, isometrically embeds into the extension graph $\Gamma^{e}$ (see Corollary 3.29). (Here, vertex groups are not very relevant. For example, the extension graph for infinite countable vertex groups is isomorphic to the extension graph of a RAAG with the same defining graph (see [Gen17, Fact 8.25]).) However, this fact for RAAGs means that when the asymptotic dimension of a defining graph is 0, the asymptotic dimension of the extension graph is at most 1. We generalize this as follows.
@@
In this paper, we present one application of Theorem 1.2 to graph wreath product. In the forthcoming paper, we present more applications to analytic properties of graph product. Given a group $G$ acting on a simplicial graph $\Gamma$ and another group $H$, we can assign vertex groups $\mathcal{G}=\{G_{v}\}_{v\in V(\Gamma)}$ by setting $G_{v}=H$ for every $v\in V(\Gamma)$, and the group $G$ acts on the graph product $\Gamma\mathcal{G}$ as group automorphisms by permuting vertex groups according to the action $G\curvearrowright\Gamma$. This action $G\to\mathrm{Aut}(\Gamma\mathcal{G})$ defines the semi-direct product $\Gamma\mathcal{G}\rtimes G$, which was called graph wreath product in [KM16]. Interestingly, we can consider this construction as interpolation of wreath product and free product. Indeed, given groups $G$ and $H$, the group $G$ acts on the set $G$ by left multiplication. If $\Gamma$ is the complete graph with $V(\Gamma)=G$, then we have $\Gamma\mathcal{G}\rtimes G=H\wr G$. If $\Gamma$ is the graph with $V(\Gamma)=G$ having no edge, then we have $\Gamma\mathcal{G}\rtimes G=H\ast G\,(=\ast_{g\in G}gHg^{-1}\rtimes G)$.
In [KM16], Kropholler and Martino studied homotopical finiteness conditions, i.e. type $\mathrm{F}_{n}$, of graph wreath product. In Theorem 1.3 and Corollary 1.4 below, we prove geometric properties of graph wreath product, which provide a new construction of relatively hyperbolic groups and hyperbolic groups.
@@
    For any $v\in V(\Gamma)$, $\mathrm{Stab}_{G}(v)$ is finitely generated.
@@
It seems interesting to investigate properties of hyperbolic groups constructed in Corollary 1.4. For example, studying residual finiteness of these hyperbolic groups might shed light on the long-standing open problem asking whether every hyperbolic group is residually finite, because non residually finite groups can be easily constructed by wreath product. Indeed, wreath product of two groups $H\wr G$ is residually finite if and only if $H$ and $G$ are residually finite, and either $H$ is abelian or $G$ is finite (see [Gru57, Theorem 3.2]). Even if the hyperbolic groups in Corollary 1.4 are residually finite, it is still interesting to investigate where this transition occurs as graph wreath product changes from wreath product to free product.
This paper is organized as follows. In Section 2, we explain preliminary definitions and known results that are necessary in this paper. In Section 3, we define the extension graph of graph product and study its geometry. In Section 4, we study asymptotic dimension of the extension graph and prove Theorem 1.1 by introducing an auxiliary graph that we call the coned-off extension graph. In Section 5, we discuss how properties of a defining graph carry over to the extension graph and prove Theorem 1.2 by using results in Section 3. In Section 6, we discusses the application to graph wreath product and prove Theorem 1.3 and Corollary 1.4.
Acknowledgment. I would like to thank Anthony Genevois and Denis Osin for helpful discussions. I would also like to thank Peter Kropholler for helpful comments on an earlier draft. I would also like to thank an anonymous referee for suggesting Corollary 3.16 (1) and for helpful comments, which improved exposition of the paper.
@@ 2 Preliminaries
$\mathrm{Lk}_{X}(x)$ $=\{y\in V(X)\mid(x,y)\in E(X)\},$
$\mathrm{St}_{X}(x)$ $=\{x\}\cup\mathrm{Lk}_{X}(x),$
$\mathrm{Elk}_{X}(x)$ $=\{(x,y)\in E(X)\mid y\in\mathrm{Lk}_{X}(x)\}.$
@@
Definition 2.5.
Let $(X,d_{X})$ be a metric space. For a subset $A\subset X$, the diameter of AA $\mathrm{diam}_{X}(A)\in[0,\infty]$ is defined by $\mathrm{diam}_{X}(A)=\sup_{x,y\in A}d_{X}(x,y)$. For $A\subset X$ and $r\in\mathbb{R}_{>0}$, we define $\mathcal{N}_{X}(A,r)\subset X$ by $\mathcal{N}_{X}(A,r)=\{y\in X\mid\exists\,x\in A,d_{X}(x,y)\leq r\}$. When $A$ is a singleton i.e. $A=\{x\}$ with $x\in X$, we denote $\mathcal{N}_{X}(\{x\},r)$ by $\mathcal{N}(x,r)$ for brevity, that is, $\mathcal{N}(x,r)=\{y\in X\mid d_{X}(x,y)\leq r\}$. For two subsets $A,B\subset X$, we define $d_{X}(A,B)\in\mathbb{R}_{\geq 0}$ by $d_{X}(A,B)=\inf_{x\in A,y\in B}d_{X}(x,y)$.
@@ 2.1 Graph products of groups
Definition 2.7.
Let $G$ be a group. For $g,h\in G$, we define $g,h]\in G$, by $g,h]=ghg^{-1}h^{-1}$. For subsets $A,B\subset G$, we define $AB,[A,B]\subset G$ by $AB=\{gh\in G\mid g\in A,h\in B\}$ and $A,B]=\{[g,h]\in G\mid g\in A,h\in B\}$. For a subset $A\subset G$, we denote by $\langle A\rangle$ the subgroup of $G$ generated by $A$ and also by $\langle\!\langle A\rangle\!\rangle$ the normal subgroup of $G$ generated by $A$, that is, $\langle\!\langle A\rangle\!\rangle=\langle\bigcup_{g\in G}gAg^{-1}\rangle$.
@@
Let $\Gamma$ be a simplicial graph and $\mathcal{G}=\{G_{v}\}_{v\in V(\Gamma)}$ be a collection of groups. For any $g\in\Gamma\mathcal{G}$ with $\neq 1, g=g$$g_{1}\cdots g_{n}$ is a normal form of $g$ if and only if for any pair $(i,j)$ with $1\leq i<j\leq n$ satisfying $v_{i}=v_{j}$, there exists $k$ with $i<k<j$ such that $(v_{k},v_{i})\notin E(\Gamma)$. Also, if $g=g_{1}\cdots g_{n}$ and $g=h_{1}\cdots h_{m}$ are normal forms of $g$, then $n=m$ and we obtain one from the other by finite steps of syllable shuffling.
@@ 2.2 The crossing graph and the contact graph of a quasi-median graph
Let $X$ be a quasi-median graph. A hyperplane is an equivalence class of edges of $X$, where two edges $e$ and $f$ of $X$ are defined to be equivalent if there exists a sequence of edges $e=e_{0},\cdots,e_{n}=f$ of $X$ such that, for every $1\leq i\leq n-1$, $e_{i}$ and $e_{i+1}$ are either two sides of a triangle or opposite sides of a square. We denote by $\mathcal{H}(X)$ the set of all hyperplanes of $X$. For a hyperplane $J$ of $X$, we denote by $N^{\prime}(J)$ the set of the endpoints of all the edges in $J$. A carrier of $J\in\mathcal{H}(X)$, denoted by $N(J)$, is the induced subgraph on $N^{\prime}(J)$. The crossing graph of $X$, denoted by $\Delta X$, is a graph with $V(\Delta X)=\mathcal{H}(X)$, where two distinct vertices $J_{1},J_{2}\in\mathcal{H}(X)$ are adjacent if there exists a square in $X$ that contains both an edge in $J_{1}$ and an edge in $J_{2}$. The contact graph of $X$, denoted by $\mathcal{C}X$, is a graph with $V(\mathcal{C}X)=\mathcal{H}(X)$, where two distinct vertices $J_{1},J_{2}\in\mathcal{H}(X)$ are adjacent if $N^{\prime}(J_{1})\cap N^{\prime}(J_{2})\neq\emptyset$.
@@ 3.1 Definition and basic properties of the extension graph of graph product
In this section, we record basic properties of the extension graph. Some of the lemmas here are already available or can be easily deduced from the literature (e.g. [AM15]).
@@
$V(\Gamma^{e})$ $=\{gG_{v}g^{-1}\in 2^{\Gamma\mathcal{G}}\mid v\in V(\Gamma),g\in\Gamma\mathcal{G}\},$
$E(\Gamma^{e})$ $=\{(gG_{v}g^{-1},hG_{w}h^{-1})\in V(\Gamma^{e})^{2}\mid\text{$gG_{v}g^{-1}\neq hG_{w}h^{-1}$ and $[gG_{v}g^{-1},hG_{w}h^{-1}]=\{1\}$}\}.$
@@
We can assume $h=1$ without loss of generality, because we can consider $(h^{-1}g.G_{v},G_{w})\in E(\Gamma^{e})$ in general case. Since $G_{v}$ and $G_{w}$ are non-trivial and we have $(g.G_{v},G_{w})\in E(\Gamma)$, there exist $a\in G_{v}\setminus\{1\}$ and $b\in G_{w}\setminus\{1\}$ such that $gag^{-1},b]=1$. By $(gag^{-1})b(gag^{-1})^{-1}=b$ and Corollary 3.5, we have $\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)$. This and Lemma 3.4 imply $v\in\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)$. Suppose $v=w$ for contradiction, then we have $\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)=\mathrm{St}_{\Gamma}(v)$. This implies $\mathrm{supp}(g)\subset\mathrm{St}_{\Gamma}(v)$ by Lemma 3.3 and Lemma 3.4. Hence, $g.G_{v}=G_{v}=G_{w}$. This contradicts $g.G_{v}\neq G_{w}$ by $(g.G_{v},G_{w})\in E(\Gamma^{e})$. Hence, $v\neq w$. By this and $v\in\mathrm{St}_{\Gamma}(w)$, we have $(v,w)\in E(\Gamma)$. By Lemma 3.4, there exist $h_{1},h_{2},h_{3}\in\Gamma\mathcal{G}$ such that $g=h_{1}h_{2}h_{3}$, $\mathrm{supp}(h_{1})\cup\{v\}=\mathrm{supp}(gag^{-1})$, $\mathrm{supp}(h_{2})\subset\mathrm{Lk}_{\Gamma}(v)$, and $\mathrm{supp}(h_{3})\subset\{v\}$. By $\mathrm{supp}(h_{2}h_{3})\subset\mathrm{supp}(h_{2})\cup\mathrm{supp}(h_{3})\subset\mathrm{St}_{\Gamma}(v)$, we have $g.G_{v}=h_{1}.G_{v}$. On the other hand, by $\mathrm{supp}(h_{1})\subset\mathrm{supp}(gag^{-1})\subset\mathrm{St}_{\Gamma}(w)$, we have $G_{w}=h_{1}.G_{w}$. ∎
@@
(1) Let $x\in V(\Gamma)$ and $g\in\Gamma\mathcal{G}$ satisfy $(v,g.x)\in E(\Gamma^{e})$. By Lemma 3.8 and $\mathrm{Lk}_{\Gamma}(v)=\{w\}$, we have $x=w$ and there exists $k\in\Gamma\mathcal{G}$ such that $(k.v,k.w)=(v,g.x)$. This implies $k\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v)$. Hence, we have $g.x=k.w=w$ since we have $k\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v)=G_{v}\times G_{w}$ by Lemma 3.6. Thus, $\mathrm{Lk}_{\Gamma^{e}}(v)=\{w\}$.
@@
    For any $a,b\in V(\Gamma^{e})$, we have $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)=\langle H\mid H\in\mathrm{St}_{\Gamma^{e}}(a)\cap\mathrm{St}_{\Gamma^{e}}(b)\rangle$. In particular, $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\neq\{1\}$ if and only if $d_{\Gamma^{e}}(a,b)\leq 2$.
@@
    If, in addition, the collection $\mathcal{G}=\{G_{v}\}_{v\in V(\Gamma)}$ is non-trivial finite groups, then for any $v,w\in V(\Gamma)$ with $v\neq w$ and any $g\in\Gamma\mathcal{G}$, the group $\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}$ is finite.